# Leveraging Multi-source and Heterogeneous Signals for Fatigue Detection

### v2

@@ II-C Open Fatigue Detection Datasets
A review of relevant literature has identified several open datasets for fatigue detection. As shown in Table I, the use of diverse physiological signals reaffirms our observation that researchers tend to select those sensors that are readily accessible and convenient for their studies. Meanwhile, we also note that certain signals are commonly used across diverse datasets, including HR, PPG, EDA, Acceleration(ACC) and ST.
The commonality and diversity across datasets present both challenges and opportunities for fatigue detection. For example, while FatigueSet [51], MEFAR [52], and VPFD [40] share similar types of data, they use different sensors as detailed in a later section - Empatica E4 for the first two datasets and Google Pixel Watch 2 for the third. These variations in data collection methods, sensor specifications, and pre-processing techniques can hinder standardization and generalization.
FatigueSet [51] and MEFAR [52], collected under controlled laboratory conditions, include high-fidelity signals—such as EEG in both datasets and ECG in FatigueSet—that are often impractical for real-world deployment. In contrast, the VPFD [40] dataset was designed for real-world use and thus relies on more accessible sensors except for EEG and ECG. This distinction directly motivates our evaluation of the central hypothesis posed in the Introduction: Can fatigue-relevant knowledge learned from high-resolution EEG and ECG sources be effectively transferred to enhance fatigue prediction in settings like VPFD, where only context-appropriate sensors are available? In the following sections, we present the results of this exploration.
@@
MePhy [55] 60 8
OperEYEVP [56] 10 10
WESAD [57] 15 13
FatigueSet [51] 12 13
MEFAR [52] 23 28
@@
III Heterogeneous and Multi-source Fatigue Detection Framework
@@ III-A Problem Definition
  • A target domain dataset $\mathcal{D}_{T}=\{(x_{T}^{(i)},y_{T}^{(i)})\}_{i=1}^{n_{T}}$ with limited samples, where $x_{T}^{(i)}\in\mathbb{R}^{d_{T}}$ represents the sensor feature vector defined over $\mathcal{X}_{T}$ and $y_{T}^{(i)}\in\mathcal{Y}$ is the label.$n_{T}$ is the number of target samples and $d_{T}$ is the number of feature dimensions, i.e., number of sensors.
  • A collection of source domains $\{\mathcal{D}_{s}\}_{s=1}^{S}$ , where each $\mathcal{D}_{s}=\{(x_{s}^{(i)},y_{s}^{(i)})\}_{i=1}^{n_{s}}$ contains samples from a different sensor configuration or environment. Each $x_{s}^{(i)}\in\mathbb{R}^{d_{s}}$ defined over $\mathcal{X}_{s}$ and $y_{s}^{(i)}\in\mathcal{Y}$ $S$ is the number of source domains, $n_{s}$ is the number of target source samples and $d_{s}$ is the number of feature dimensions of $\mathcal{D}_{s}$
@@
  • Feature Space Heterogeneity: The feature spaces across different domains are partially but not completely overlapping, meaning that the set of available features (sensors) in the target domain $\mathcal{D}_{T}$ and source domains $\{\mathcal{D}_{s}\}_{s=1}^{S}$ share some common modalities, but also contain domain-specific ones. Formally, $\mathcal{X}_{T}\neq\mathcal{X}_{s}$ and $\mathcal{X}_{T}\cap\mathcal{X}_{s}\neq\emptyset$ for $s=1,\ldots,S$ Note that scenarios where the feature spaces are entirely disjoint are not considered in this study.
  • Data Distribution Heterogeneity: The data distributions across different domains are not identical, meaning that the statistical properties of the data in the target domain $\mathcal{D}_{T}$ and source domains $\{\mathcal{D}_{s}\}_{s=1}^{S}$ are different. Formally, $P_{T}\neq P_{s}$ for $s=1,\ldots,S$ , where $P_{T}$ and $P_{s}$ represent the data distributions of the target and source domains, respectively.
@@
  • $\mathcal{X}^{s-}:=\mathcal{X}_{s}\setminus\mathcal{X}_{T}^{\cap}$
@@
  • $\mathcal{X}^{s+}:=\mathcal{X}_{T}\setminus\mathcal{X}_{s}^{\cap}$
@@
Here, we assume that each sensor, regardless of its type or modality, is available in at least one domain. Besides, we also have $\mathbb{P}_{T}^{\cap}\neq\mathbb{P}_{s}^{\cap}$, where $\mathbb{P}_{T}^{\cap}$ and $\mathbb{P}_{s}^{\cap}$ is the distribution over $\mathcal{X}^{\cap}$ of $\mathcal{D}_{T}$ and $\mathcal{D}_{s}$, respectively. Here, our task is to enhance the performance of a fatigue detector $f$ in the target domain by leveraging data from multiple heterogeneous source domains. Specifically, we aim to improve the fatigue detection accuracy in the target domain by effectively integrating and aligning data from various source domains, despite their heterogeneous feature spaces and data distributions.
@@ III-B Theoretical Analysis
To gain a deeper understanding of the problem and our proposed solution, we theoretically analyze the generalization bound associated with it. For simplicity and ease of understanding, in our theoretical analysis, we only considered one source domain, which can readily extends to multiple source domains without additional conceptual difficulty. For a random variable $x$ generated from a distribution $\mathbb{P}$, we use $\mathbb{E}_{x\sim\mathbb{P}}$ to denote the expectation taken over $x$ with distribution $\mathbb{P}$. The expected generalization error of the Fatigue Detector $f(x)$ is $\mathcal{E}_{\mathbb{P}}\left(f\right)$, while the empirical generalization error is the ^$\mathcal{E}_{\hat{\mathbb{P}}}\left(f\right)$.
@@
[58] Suppose the loss function $\mathcal{L}(f(x),y)$ is R-sub-Gaussian under $x\sim\mathbb{P}$ for all $y\in\mathcal{Y}$, then, we have:
@@
Assume that $x,a$ are independent of each other,$x\sim\mathbb{P},a\sim\mathbb{Q}_{a}$, while both being conditionally dependent on variable of $y$ , Let the composite observation be defined as $x^{+}=[x,s],x^{+}\sim\mathbb{P}_{+}$, with the incorporating the variable $a$ , the $\mathcal{E}_{\mathbb{P}_{+}}$ have a tighten upper bound than $\mathcal{E}_{\mathbb{P}}$, the gap $G$ is:
@@
Given a random variable $\hat{a}$ generated from a distribution $\mathbb{Q}_{\hat{a}}$, and $\hat{x}^{+}=[x,\hat{a}]\sim\mathbb{Q}$ , then we have:
where ^$\mathcal{E}_{\hat{\mathbb{Q}}}\left(f\right)$ is the empirical generalization error over distribution $\mathbb{Q}$, $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$ is the $\mathcal{H}\Delta\mathcal{H}$ distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$. The proof is given in the Appendix.
Through Theorem 1 and Theorem 2, it is evident that without additional sensors $a$, if the distribution distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$, is less than a constant, i.e., $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})\leq\left|\sqrt{\frac{2R^{2}}{n}}\left(\sqrt{I(a,y)}-\sqrt{I(x,y)}\right)-\Delta-\epsilon_{ideal}\right|$, we can obtain a tighter generalization error bound, thereby improving the generalization performance of the model. This result validates that generating advanced sensors can enhance the performance of fatigue detection.
Thus, to effectively enhance model performance, we are motivated to minimize the distributional discrepancy $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$ between generated sensor $\hat{a}$ and potential real sensor $a$ by training a regression model $g_{s}:\mathcal{X}^{\cap}\xrightarrow{}\mathcal{X}^{s+}$ over $\mathcal{D}_{s}$. Since the $\mathcal{H}\Delta\mathcal{H}$ distance is symmetric, obeys the triangle inequality, and is bounded, we can regard it as a loss function $\ell$. Following the current theoretical work, [59]. We easily have the following results.
@@
where $R_{S}^{\ell}(g_{s},g_{S}^{*})=\mathbb{E}_{x\sim\mathbb{P}_{s}}\ell(g_{s}(x),g_{s}^{*}(x))$ and $\epsilon=R_{T}^{\ell}(g_{T}^{*})+R_{S}^{\ell}(g_{T}^{*},g_{s}^{*})$. Since $\hat{a}$ is generated by $g_{s}$. The $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$ is bounded by $R_{S}^{\ell}(g_{s},g_{S}^{*})$ and $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$ Following this idea, to reduce the distributional discrepancy $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$, we also require to reduce the distribution gap $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$.
@@ III-C Learning Objective
Motivated by Theorem 1 and 2, we provide our heterogeneous and multi-source framework. Specifically, given a target domain $\mathcal{D}_{T}$, our learning objective is to detect fatigue with knowledge transferring from multiple source domains (e.g., high-fidelity knowledge extracted from controlled environments) $\{\mathcal{D}_{s}\}$ for $s=1,\ldots,S$. Here, as a naive attempt, we aim to transfer knowledge by utilizing the remaining domains to generate the missing sensors (i.e., $\mathcal{X}^{s+}$) in $\mathcal{D}_{s}$ The learning objective is defined as follows:
$\mathcal{L}=\min_{g_{s},f}\underbrace{\sum_{s=1}^{S}\mathbb{E}_{\mathcal{D}_{s}}[\mathcal{L}_{rec}(g_{s}(\mathcal{X}^{\cap},\mathcal{X}^{s+})}_{\text{SensorReconstruction}}+\lambda\underbrace{\mathbb{E}_{\mathcal{D}_{T}}[\mathcal{L}_{task}(f(\tilde{\mathcal{X}_{T}}),y)]}_{\text{Task-Aware Detection}}$ (5)
where $g_{s}$ is the feature generator for generating $\mathcal{X}^{s+}$, $f$ is the classifier function for fatigue detection . $\mathcal{L}_{rec}$ represents the reconstruction loss (e.g., MSE for regression). $\mathcal{L}_{task}$ denotes the task-specific loss (e.g., MSE for regression, CE for classification), $\lambda$ is the adaptive coefficients from uncertainty weighting. $\tilde{\mathcal{X}_{T}}$ represents the reconstructed data from domain $\mathcal{D}_{T}$ This data has been enhanced or completed using the information generated by the feature generator $g_{s}$ from other domains.
Moreover, motivated by Theorem 3 , a batch normalization step [60], referred to as BatchNorm, is applied before fusion, to reduce $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$. Additionally, to optimize the objective function in Equation 5, particularly the task-specific loss, we adopt the Jacobian Norm Regularization method [61]-referred to as JacobianNorm later.
@@ III-D Framework
The overall framework is summarized in Figure 1, with the detailed procedures outlined in Algorithms 1 and 2. Specifically, we propose a proof-of-concept approach that enables cross-domain knowledge integration for improved fatigue detection in heterogeneous sensor settings. Specifically, the method is designed to handle data collected from different sensor configurations with partially overlapping feature spaces. When new data from a different sensor setup becomes available, the framework compares its feature space to that of the target domain. If discrepancies are found—i.e., missing modalities—the method initiates a targeted imputation process to infer the missing signals. The aligned and augmented data are then incorporated into the model to enhance prediction accuracy in the target domain.
Algorithm 1 Multi-source Fatigue Detection
@@
1:  Compute modality-difference sets
@@
Figure 1: Overall framework

### v3

@@ II-C Open Fatigue Detection Datasets
A review of relevant literature has identified several open datasets for fatigue detection. As shown in Table I, the use of diverse physiological signals reaffirms that researchers tend to select sensors that are readily accessible and convenient. Meanwhile, certain signals recur across datasets, including HR, PPG, EDA, acceleration (ACC), and ST.
The commonality and diversity across datasets present both challenges and opportunities. For example, while FatigueSet [10.1007/978-3-030-99194-4_14], MEFAR [DERDIYOK2024109896], and VPFD [10799435] share similar data types, they use different sensors—Empatica E4 for the first two datasets and Google Pixel Watch 2 for the third. These variations in data collection methods, sensor specifications, and pre-processing can hinder standardization and generalization.
In addition to these sensor-level discrepancies, label heterogeneity also presents a major challenge for joint modeling and cross-dataset transfer. For instance, some datasets use continuous fatigue scores, whereas others rely on binary labels such as “fatigued” or “not fatigued.” Even among continuous labels, threshold definitions may differ (e.g., fatigue defined as scores >3>3 or >5>5 on a subjective scale). These inconsistencies need to be addressed to enable more consistent model transfer across datasets.
FatigueSet [10.1007/978-3-030-99194-4_14] and MEFAR [DERDIYOK2024109896], collected under controlled laboratory conditions, include high-fidelity signals—such as EEG in both datasets and ECG in FatigueSet—that are often impractical for real-world deployment. In contrast, the VPFD [10799435] dataset was designed for real-world use and thus relies on more accessible sensors except for EEG and ECG. This distinction directly motivates our evaluation of the central hypothesis posed in the Introduction: Can fatigue-relevant knowledge learned from high-resolution EEG and ECG sources be effectively transferred to enhance fatigue prediction in settings like VPFD, where only context-appropriate sensors are available? Note that the three datasets provide a relatively homogeneous foundation for our evaluation, as all rely on self-reported binary fatigue scores. The only distinction lies in FatigueSet, which separates mental and physical fatigue, whereas MEFAR and VPFD do not. For consistency across datasets, we combined both mental and physical fatigue into a unified label. In the following sections, we present the results of this exploration.
@@
MePhy [Gabbi2024UnderstandingFT] 60 8
OperEYEVP [s23136197] 10 10
WESAD [10.1145/3242969.3242985] 15 13
FatigueSet [10.1007/978-3-030-99194-4_14] 12 13
MEFAR [DERDIYOK2024109896] 23 28
@@
III Heterogeneous and Multi source Fatigue Detection Framework
@@ III-A Problem Definition
[…]
  • A collection of source domains $\{\mathcal{D}_{s}\}_{s=1}^{S}$, where each $\mathcal{D}_{s}=\{(x_{s}^{(i)},y_{s}^{(i)})\}_{i=1}^{n_{s}}$ contains samples from a different sensor configuration or environment. Each $x_{s}^{(i)}\in\mathbb{R}^{d_{s}}$ is defined over $\mathcal{X}_{s}$ and $y_{s}^{(i)}\in\mathcal{Y}$. $S$ is the number of source domains, $n_{s}$ is the number of samples in source domain $s$, and $d_{s}$ is the feature dimensionality of $\mathcal{D}_{s}$.
@@
  • Feature Space Heterogeneity: For each source $s$, $\mathcal{X}_{T}\neq\mathcal{X}_{s}$ may hold while $\mathcal{X}_{T}\cap\mathcal{X}_{s}\neq\emptyset$; when $\mathcal{X}_{T}=\mathcal{X}_{s}$, our setting reduces to an identical-modality special case (no imputation required).
  • Data Distribution Heterogeneity: $\mathbb{P}_{T}\neq\mathbb{P}_{s}$ for $s=1,\ldots,S$, where $\mathbb{P}_{T}$ and $\mathbb{P}_{s}$ denote the distributions of target and source domains, respectively.
@@
  • $\mathcal{X}^{s-}:=\mathcal{X}_{T}\setminus\mathcal{X}^{\cap}$
@@
  • $\mathcal{X}^{s+}:=\mathcal{X}_{s}\setminus\mathcal{X}^{\cap}$
@@
Here, we assume that each sensor, regardless of its type or modality, is available in at least one domain. Besides, we also have $\mathbb{P}_{T}^{\cap}\neq\mathbb{P}_{s}^{\cap}$, where $\mathbb{P}_{T}^{\cap}$ and $\mathbb{P}_{s}^{\cap}$ is the distribution over $\mathcal{X}^{\cap}$ of $\mathcal{D}_{T}$ and $\mathcal{D}_{s}$, respectively. Here, our task is to enhance the performance of a fatigue detector $f$ in the target domain by leveraging data from multiple heterogeneous source domains. Specifically, under distribution shift, we seek to generate the features missing in the target domain by leveraging multiple heterogeneous source datasets, thereby enhancing the model’s fatigue-detection performance. For completeness, for each source $s$ we learn a modality-imputation mapping $g_{s}:\mathcal{X}^{\cap}\to\mathcal{X}^{s+}$ on $\mathcal{D}_{s}$, construct $\hat{a}_{T}^{(s)}:=g_{s}(x^{\cap})$, and form the augmented target feature $\tilde{x}_{T}^{(s)}=[x^{\cap},\hat{a}_{T}^{(s)}]$. In the multi-source case, we concatenate or fuse the reconstructed features $\{\hat{a}_{T}^{(s)}\}_{s=1}^{S}$ to obtain the final augmented representation.
@@ III-B Theoretical Analysis
To gain a deeper understanding of the problem and our proposed solution, we theoretically analyze the generalization bound associated with it. For simplicity and ease of understanding, in our theoretical analysis, we only considered one source domain, which can readily extends to multiple source domains without additional conceptual difficulty. For a random variable $x$ generated from a distribution $\mathbb{P}$, we use $\mathbb{E}_{x\sim\mathbb{P}}$ to denote the expectation taken over $x$ with distribution $\mathbb{P}$. The expected generalization error of the Fatigue Detector $f(x)$ is $\mathcal{E}_{\mathbb{P}}\left(f\right)$, while the empirical generalization error is the $\mathcal{E}_{\hat{\mathbb{P}}}\left(f\right)$. Throughout this section, $I(x;y)$ denotes mutual information between the labels $y$ and the training features $x$, i.e., sensors in our task.
@@
[bu_tightening_2019] Suppose the loss function $\mathcal{L}(f(x),y)$ is R-sub-Gaussian under $x\sim\mathbb{P}$ for all $y\in\mathcal{Y}$, then, we have:
$\mathcal{E}_{\mathbb{P}}\left(f\right)\leq\mathcal{E}_{\hat{\mathbb{P}}}\left(f\right)+\sqrt{\frac{2R^{2}}{n}\,{\color[rgb]{0,0,0}I(x;y)}}$ (1)
Let $a$ denote a label-related advanced sensor and $\hat{a}$ its imputed counterpart. Define $x^{+}=[x,a]$ with joint distribution $\mathbb{P}_{+}$ over $(x^{+},y)$, and $\hat{x}^{+}=[x,\hat{a}$] with joint distribution $\mathbb{Q}$ over $(\hat{x}^{+},y)$.
@@
Assume that $x,a$ are independent of each other, $x\sim\mathbb{P},a\sim\mathbb{Q}_{a}$, while both being conditionally dependent on variable of $y$, Let the composite observation be defined as $^{+}=[x,a],x++,x^{+}\sim\mathbb{P}_{+}$, with the incorporating the variable $a$ , the $\mathcal{E}_{\mathbb{P}_{+}}$ have a tighten upper bound than $\mathcal{E}_{\mathbb{P}}$, the gap $G$ is:
$G=\Delta+\sqrt{\frac{2R^{2}}{n}}\left(\sqrt{{\color[rgb]{0,0,0}I(x^{+};y)}}-\sqrt{{\color[rgb]{0,0,0}I(x;y)}}\right)<0$ (2)
@@
Given a random variable $^{\color[rgb]{0,0,0}\hat{a}}$ generated from a distribution $^{\color[rgb]{0,0,0}\mathbb{Q}_{\hat{a}}}$, and $\hat{x}^{+}=[x,{\color[rgb]{0,0,0}\hat{a}}$$\sim{\color[rgb]{0,0,0}\mathbb{Q}}$ , then we have:
$\mathcal{E}_{\mathbb{P}_{+}}\left(f\right)\leq\mathcal{E}_{\hat{\mathbb{Q}}}\left(f\right)+\epsilon_{ideal}+{\color[rgb]{0,0,0}d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})}+\sqrt{\frac{2R^{2}}{n}{\color[rgb]{0,0,0}\,I(x;y)}}$ (3)
[…]
Through Theorem 1 and Theorem 2, it is evident that without additional sensors $a$, if the distribution distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$, is less than a constant, i.e., $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})\leq\left|\sqrt{\frac{2R^{2}}{n}}\left(\sqrt{{\color[rgb]{0,0,0}I(x^{+};y)}}-\sqrt{{\color[rgb]{0,0,0}I(x;y)}}\right)-\Delta-\epsilon_{ideal}\right|$, we can obtain a tighter generalization error bound, thereby improving the generalization performance of the model. This result validates that generating advanced sensors can enhance the performance of fatigue detection.
[…]
@@
[…]
@@ III-C Learning Objective
Motivated by Theorem 1 and 2, we provide our heterogeneous and multi source framework. Specifically, given a target domain $\mathcal{D}_{T}$, our learning objective is to detect fatigue with knowledge transferring from multiple source domains (e.g., high fidelity knowledge extracted from controlled environments) $\{\mathcal{D}_{s}\}$ for $s=1,\ldots,S$. Here, as a naive attempt, we aim to transfer knowledge by utilizing the remaining domains to generate the missing sensors (i.e., $\mathcal{X}^{s+}$) in $\mathcal{D}_{s}$ . The learning objective is defined as follows:
$\mathcal{L}=\min_{g_{s},f}$ $\underbrace{\sum_{s=1}^{S}\mathbb{E}_{\mathcal{D}_{s}}[\mathcal{L}_{rec}(g_{s}(\mathcal{X}^{\cap},\mathcal{X}^{s+})}_{\text{Sensor Reconstruction}}+\underbrace{[\mathcal{L}_{dis}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap}))]}_{\text{Domain Alignment}}$ (5)
$+$ $\underbrace{\mathbb{E}_{\mathcal{D}_{T}}[\mathcal{L}_{task}(f(\tilde{\mathcal{X}_{T}}),y)]}_{\text{Task Aware Detection}}$
where $g_{s}$ is the feature generator for generating $\mathcal{X}^{s+}$, $f$ is the classifier function for fatigue detection . $\mathcal{L}_{rec}$ represents the reconstruction loss (e.g., MSE for regression). $\mathcal{L}_{task}$ denotes the task specific loss (e.g., MSE for regression, CE for classification), $\tilde{\mathcal{X}_{T}}$ represents the reconstructed data from domain $\mathcal{D}_{T}$ . This data has been enhanced or completed using the information generated by the feature generator $g_{s}$ from other domains.
Moreover, motivated by Theorem 3, we introduce a domain alignment term to reduce $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{T}^{\cap},\mathbb{P}_{s}^{\cap})$. Specifically, a batch normalization step[ioffe2015batchnormalizationacceleratingdeep], referred to as BatchNorm, is applied before fusion. Additionally, to optimize the objective function in Equation 5, particularly the task specific loss, we adopt the Jacobian Norm Regularization method [Li_2023] referred to as JacobianNorm later.
@@ III-D Framework
For the problem defined above, we propose a framework that enables cross-domain knowledge integration to improve fatigue detection under heterogeneous sensor configurations. The overall framework is given in Fig.1, with the detailed procedures outlined in Algorithm 1. The method is designed to handle data collected from different sensor configurations with partially overlapping feature spaces. When new data from a different sensor setup becomes available, the framework compares its feature space to that of the target domain. If discrepancies are found—i.e., missing modalities—the method initiates a targeted imputation process to infer the missing signals. This imputation process is detailed in Algorithm 2 and illustrated in Fig.2. The aligned and augmented data is then incorporated into the model to enhance prediction accuracy in the target domain as illustrated in Fig.3.
Algorithm 1 Multi source Fatigue Detection
@@
1: Compute modality difference sets
@@
Figure 1: Overall framework.
Figure 2: Sensor Imputation Module.
Figure 3: Fatigue Detection model.